OpenStudy (anonymous):

5^-8 x 5^4 x= multiply not a variable

OpenStudy (anonymous):

so am i trying to find x?

OpenStudy (anonymous):

==5^4/5^8=1/5^4

OpenStudy (anonymous):

no you are mltiplying them . . . i need it in a fraction. .

OpenStudy (anonymous):

do i know you mimi?

OpenStudy (anonymous):

idk do you?

OpenStudy (agreene):

\[5^{-8}\times 5^4= \frac{5^4}{5^{-8}}=\frac{1}{5^4}=\frac{1}{625}=0.0016\]

OpenStudy (anonymous):

5^-8 x 5^4 using the law a^m x a^n = a^m+n 5^(-8)+4 = 5^-4 using the law a^-m = 1/a^m 1 1 5^-4 = ----- = ----- 5^4 625

OpenStudy (agreene):

in that first fraction there shouldnt be a -8, just an 8.

OpenStudy (anonymous):

im gonna write this all down =)

OpenStudy (anonymous):

knowledge of exponential laws is very important for solving such sums...☺

OpenStudy (anonymous):

ohhh no no, its just that my cousin uses the same name, lol sorry

OpenStudy (anonymous):

hahaha thats ok =)

OpenStudy (anonymous):

are you korean? or do u listen to kpop?